### v1

Abstract
In this paper, we study the complexity of the periodic temporal graph realization problem with respect to upper bounds on the fastest path durations among its vertices. This constraint with respect to upper bounds appears naturally in transportation network design applications where, for example, a road network is given, and the goal is to appropriately schedule periodic travel routes, while not exceeding some desired upper bounds on the travel times. This approach is in contrast to verification applications of the graph realization problems, where exact values for the distances (respectively, fastest travel times) are given, following some kind of precise measurement. In our work, we focus only on underlying tree topologies, which are fundamental in many transportation network applications.
As it turns out, the periodic upper-bounded temporal tree realization problem (TTR) has a very different computational complexity behavior than both (i) the classic graph realization problem with respect to shortest path distances in static graphs and (ii) the periodic temporal graph realization problem with exact given fastest travel times (which was recently introduced). First, we prove that, surprisingly, TTR is NP-hard, even for a constant period $\Delta$ and when the input tree $G$ satisfies at least one of the following conditions: (a) $G$ has a constant diameter, or (b) $G$ has constant maximum degree. In contrast, when we are given exact values of the fastest travel delays, the problem is known to be solvable in polynomial time. Second, we prove that TTR is fixed-parameter tractable (FPT) with respect to the number of leaves in the input tree $G$, via a novel combination of techniques for totally unimodular matrices and mixed integer linear programming.
keywords:
Temporal graph, periodic temporal labeling, fastest temporal path, graph realization, temporal connectivity.
category:
[…]
1 Introduction
A temporal (or dynamic) graph is a graph whose topology is subject to discrete changes over time. This paradigm reflects the structure and operation of a great variety of modern networks; social networks, wired or wireless networks whose links change dynamically, transportation networks, and several physical systems are only a few examples of networks that change over time [54, 58, 42]. Inspired by the foundational work of Kempe et al. [43], we adopt here a simple model for temporal graphs, in which the vertex set remains unchanged while each edge is equipped with a set of integer time labels.
Definition 1.1 (temporal graph [43]).
A temporal graph is a pair $(G,\lambda)$, where $G=(V,E)$ is an underlying (static) graph and $\lambda:E\rightarrow 2^{\mathbb{N}}$ is a labeling function which assigns to every edge of $G$ a finite set of discrete time labels.
Here, whenever $t\in\lambda(e)$, we say that the edge $e$ is active or available at time $t$. In the context of temporal graphs, where the notion of vertex adjacency is time-dependent, the notions of path and distance also need to be redefined. The most natural temporal analogue of a path is that of a temporal (or time-dependent) path, which is a path of the underlying static graph whose time labels strictly increase along the edges of the path. The duration of a temporal path is the number of discrete time steps needed to traverse it. Finally, a temporal path from vertex $u$ to vertex $v$ is fastest if it has the smallest duration among all temporal paths from $u$ to $v$.
The graph realization problem with respect to some graph property $\mathcal{P}$ is to compute a graph that satisfies $\mathcal{P}$, or to decide that no such graph exists. The main motivation for graph realization problems stems both from network design and from “verification” applications in engineering. In network design applications, the goal is to design network topologies having a desired property [4, 38]. One important application domain of network design problems is that of transportation network design where, for example, a road network is given, and the goal is to appropriately schedule periodic travel routes while not exceeding some desired upper bounds on travel times. On the other hand, in verification applications, given the outcomes of some exact experimental measurements, the aim is to (re)construct an input network that complies with them. If such a reconstruction is not possible, this proves that the measurements are incorrect or implausible (resp. that the algorithm making the computations is incorrectly implemented).
One example of a graph realization problem is the recognition of probe interval graphs, in the context of the physical mapping of DNA, see [51, 50] and [36, Chapter 4]. Analyzing the computational complexity of the graph realization problems for various natural and fundamental graph properties $\mathcal{P}$ requires a deep understanding of these properties. Among the most studied parameters for graph realization are constraints on the distances between vertices [9, 8, 40, 59, 60, 62, 18, 12, 19], on the vertex degrees [35, 37, 39, 7, 25], on the eccentricities [6, 41, 11, 49], and on connectivity [34, 30, 17, 31, 32, 37], among others. Although the majority of studies of graph realization problems concern static graphs, the temporal (periodic) graph realization problem with respect to given (exact) delays of the fastest temporal paths among vertices has been recently studied in [46, 27], motivated by verification applications where exact measurements are given.
In this paper, we consider the (periodic) temporal graph realization problem with respect to the durations of the fastest temporal paths from a network design perspective, i. e., where only upper bounds on the durations of fastest temporal paths are given. Our work is motivated by transportation network applications, where an underlying tree graph structure is of fundamental importance [5, 10, 13]. For example, many airlines or railway transportation networks are even star graphs (having a big hub at the center of the network, e.g. in the capital city). The formal definition of our problem is as follows (see Section 2 for formal definitions of all used terminology).
Periodic Upper-Bounded Temporal Tree Realization (TTR)
Input: A tree $G=(V,E)$ with $V=\{v_{1},v_{2},\ldots,v_{n}\}$, an $n\times n$ matrix $D$ of positive integers, and a positive integer $\Delta$. Question: Does there exist a $\Delta$-periodic labeling $\lambda:E\rightarrow\{1,2,\ldots,\Delta\}$ such that, for every $i,j$, the duration of the fastest temporal path from $v_{i}$ to $v_{j}$ in the $\Delta$-periodic temporal graph $(G,\lambda,\Delta)$ is at most $D_{i,j}$.
Many natural and technological systems exhibit a periodic temporal behavior. This is true even more on transportation networks [3], where the goal is to build periodic schedules of transportation units (e.g. trains, buses, or airplanes). The most natural constraint, while designing such periodic transportation schedules, is that the fastest travel time (i. e., the duration of the fastest temporal path) between a specific pair of vertices does not exceed a specific desired upper bound. That is, if the travel time between $u$ and $v$ in the resulting schedule is shorter than this upper bound, the schedule is even better (and thus, still feasible). Periodic temporal graphs have also been studied in various different contexts (see [14, Class 8] and [3, 28, 57, 56]).
We focus on the most fundamental case of periodic temporal graphs, where every edge has the same period $\Delta$ and it appears exactly once within each period (while in the more general case, each edge $e$ has a period $\Delta_{e}$ and might appear more than once in each period). As it turns out, TTR has a very different computational complexity behavior than both (i) the classic graph realization problem with respect to shortest path distances in static graphs [40] and (ii) the periodic temporal graph realization problem with exact given fastest travel times [46]. We remark that very recently, the case where every edge can appear multiple times in each $\Delta$ period has been studied [27].
Our contribution.
Our results in this paper are given in two main directions. First, we prove that TTR is NP-hard, even for a constant period $\Delta$ and when the input tree $G$ satisfies at least one of the following conditions: (a) $G$ has a constant diameter, or (b) $G$ has constant maximum degree. On the one hand, this hardness comes in wide contrast to the classic (static) graph realization problem with respect to shortest path distances. Indeed, this static analogue is easily solvable in polynomial time in two steps [40]. On the other hand, the complexity of TTR is also surprisingly different than the periodic temporal graph realization problem with exact given fastest travel times. More specifically, when the input integer matrix $D$ gives the exact fastest travel times, the problem becomes solvable in polynomial time on trees [46].
Second, we prove that TTR is fixed-parameter tractable (FPT) with respect to the number of leaves in the input tree $G$. That is, long chains of vertices of degree 2 do not affect the complexity of the problem. To provide our FPT algorithm, we reduce TTR to a number Mixed Integer Linear Program (MILP) instances that is upper-bounded by a function of the number of leaves in the input tree. Furthermore, the number of integer variables in each MILP instance is also upper-bounded by a function of the number of leaves in the input tree. This allows us to use a known FPT algorithm for MILP parameterized by the number of integer variables [48, 20] to solve all MILP instances in FPT-time with respect to the number of leaves of the input tree. We prove that the TTR instance is a yes-instance if and only if at least one of the MILP instances admits a feasible solution. To this end, we use a novel combination of techniques for totally unimodular matrices and mixed integer linear programming to design the MILP instances in a way that if they admit a feasible solution, then we can assume that all variables are set to integer values.
Related work.
Since the 1960s, (static) graph realization problems have been extensively studied. We give an overview of related work in the introduction.
Several problems related to graph realization have been studied in the temporal setting. In most cases, the goal is to assign labels to (sets of) edges of a given static graph to achieve certain connectivity-related properties [45, 52, 1, 23, 46, 27]. In most of the mentioned works, the sought labeling is not periodic and the desired property is to establish temporal connectivity among all vertices [45, 52, 1] or a subset of vertices [45], or to restrict reachability [23]. Among all these works, periodic labelings and the duration of the temporal paths were only considered in [46, 27].
Numerous further problems related to temporal connectivity have been studied [53, 22, 55, 21, 26, 29, 63, 24, 15, 33]. Due to the added temporal dimension, the considered problems are typically computationally harder their non-temporal counterparts, and some of which do not even have a non-temporal counterpart. In many cases, restricting the underlying topology to trees has been considered to obtain tractability results [47, 44, 2].
Figure 1: Visualization of a $\Delta$-periodic temporal graph $(G,\lambda,\Delta)$ with $\Delta=5$ and the following $\Delta$-periodic labeling $\lambda:E\rightarrow\{1,2,\ldots,5\}$: $\lambda(\{v_{1},v_{2}\})=\lambda(\{v_{2},v_{3}\})=3$, $\lambda(\{v_{3},v_{4}\})=4$, and $\lambda(\{v_{4},v_{5}\})=1$. A fastest temporal path from $v_{1}$ to $v_{5}$ first traverses edge $\{v_{1},v_{2}\}$ at time $3$, then edge $\{v_{2},v_{3}\}$ a time $8$, then edge $\{v_{3},v_{4}\}$ at time $9$, and then edge $\{v_{4},v_{5}\}$ at time $11$, and has duration $9$.
2 Preliminaries and notation
We already introduced the most central notions and concepts. There are some additional definitions we need to present our proofs and results, which we give in the following.
An undirected graph $G=(V,E)$ consists of a set $V$ of vertices and a set $E\subseteq\binom{V}{2}$ of edges. For a graph $G$, we also denote by $V(G)$ and $E(G)$ the vertex and edge set of $G$, respectively. We denote an edge $e\in E$ between vertices $u,v\in V$ as a set $e=\{u,v\}$.
Let $G=(V,E)$ and $\Delta\in\mathbb{N}$, and let $\lambda:E\rightarrow\{1,2,\ldots,\Delta\}$ be an edge-labeling function that assigns to every edge of $G$ exactly one of the labels from $\{1,\ldots,\Delta\}$. Then we denote by $(G,\lambda,\Delta)$ the $\Delta$-periodic temporal graph $(G,L)$, where for every edge $e\in E$ we have $L(e)=\{\lambda(e)+i\Delta\mid i\geq 0\}$. In this case, we call $\lambda$ a $\Delta$-periodic labeling of $G$. When it is clear from the context, we drop $\Delta$ from the notation and we denote the ($\Delta$-periodic) temporal graph by $(G,\lambda)$.
A temporal $(s,z)$-walk (or temporal walk) of length $k$ from vertex $s=v_{0}$ to vertex $z=v_{k}$ in a $\Delta$-periodic temporal graph $(G,L)$ is a sequence $P=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$ of triples that we call transitions, such that for all $i\in[k]$ we have that $t_{i}\in L(\{v_{i-1},v_{i}\})$ and for all $i\in[k-1]$ we have that $t_{i}<t_{i+1}$. Moreover, we call $P$ a temporal $(s,z)$-path (or temporal path) of length $k$ if $v_{i}\neq v_{j}$ for all $i,j\in\{0,\ldots,k\}$ with $i\neq j$. Given a temporal path $P=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$, we denote the set of vertices of $P$ by $V(P)=\{v_{0},v_{1},\ldots,v_{k}\}$. A temporal $(s,z)$-path $P=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$ is fastest if for all temporal $(s,z)$-path $P^{\prime}=\left(\left(v^{\prime}_{i-1},v^{\prime}_{i},t^{\prime}_{i}\right)\right)_{i=1}^{k^{\prime}}$ we have that $t_{k}-t_{0}\leq t^{\prime}_{k^{\prime}}-t^{\prime}_{0}$. We say that the duration of $P$ is $d(P)=t_{k}-t_{0}+1$. Furthermore, the concept of travel delays is very important for our proofs.
Definition 2.1 (travel delays).
Let $(G,\lambda)$ be a $\Delta$-periodic temporal graph. Let $e_{1}=\{u,v\}$ and $e_{2}=\{v,w\}$ be two incident edges in $G$ with $e_{1}\cap e_{2}=v$. We define the travel delay from $u$ to $w$ at vertex $v$, denoted with $\tau_{v}^{u,w}$, as follows.
$\tau_{v}^{u,w}=\begin{cases}\lambda(e_{2})-\lambda(e_{1}),&\text{if }\lambda(e_{2})>\lambda(e_{1}),\\ \lambda(e_{2})-\lambda(e_{1})+\Delta,&\text{otherwise}.\end{cases}$
From Definition 2.1 we immediately get the following observation. {observation} Let $(G,\lambda)$ be a $\Delta$-periodic temporal graph. Let $e_{1}=\{u,v\}$ and $e_{2}=\{v,w\}$ be two incident edges in $G$ with $e_{1}\cap e_{2}=v$. Then we have the following.
  • $\tau_{v}^{u,w}=\Delta-\tau_{v}^{w,u}$ if $\lambda(e_{1})\neq\lambda(e_{2})$.
  • $\tau_{v}^{u,w}=\tau_{v}^{w,u}=\Delta$ if $\lambda(e_{1})=\lambda(e_{2})$.
Intuitively, the value of $\tau_{v}^{u,w}$ quantifies how long a fastest temporal path waits at vertex $v$ when first traversing edge $\{u,v\}$ and then edge $\{v,w\}$. Formally, we have the following very useful lemma.
Lemma 2.2.
Let $P=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$ be a fastest temporal $(s,z)$-path. Then we have
$d(P)=1+\sum_{i\in[k-1]}\tau_{v_{i}}^{v_{i-1},v_{i+1}}.$
Proof 2.3.
This statement can straightforwardly be shown by induction on $k$. For $k=1$ it follows from Definition 2.1 and the definition of the duration of a fastest temporal path. For $k>1$ we obtain the following.
$d(P)$ $=t_{k}-t_{0}+1=(t_{k-1}-t_{0}+1)+t_{k}-t_{k-1}$
$=(1+\sum_{i\in[k-2]}\tau_{v_{i}}^{v_{i-1},v_{i+1}})+t_{k}-t_{k-1}$
$=(1+\sum_{i\in[k-2]}\tau_{v_{i}}^{v_{i-1},v_{i+1}})+\tau_{v_{k-1}}^{v_{k-2},v_{k}}$
$=1+\sum_{i\in[k-1]}\tau_{v_{i}}^{v_{i-1},v_{i+1}}.$
Here, we use Definition 2.1 and the fact that $P$ is a fastest temporal path to obtain $t_{k}-t_{k-1}=\tau_{v_{k-1}}^{v_{k-2},v_{k}}$.
3 NP-hardness of TTR
In this section, we prove that TTR is NP-hard. In fact, it remains hard even in quite restrictive settings: when the input tree has constant diameter or constant maximum degree, and in addition the period $\Delta$ is also constant. We remark that the diameter can be seen as a measure of how “star-like” a tree is. In particular, on trees, the diameter upper bounds the treedepth. Formally, we show the following.
Theorem 3.1.
TTR is NP-hard even if $\Delta$ is constant, and
  • the input tree $G$ has constant diameter, or
  • the input tree $G$ has constant maximum degree.
To show Theorem 3.1 we present two polynomial-time many-one reductions from the NP-hard problem Monotone Not-All-Equal 3-Satisfiability (MonNAE3SAT) [Sch78] to TTR.
Monotone Not-All-Equal 3-Satisfiability (MonNAE3SAT)
Input: A set $X$ of Boolean variables and a set $C\subseteq\binom{X}{3}$ of clauses, each of which contains three variables. Question: Is there an assignment to the variables $X$ such that for each clause in $C$ at least one variable is set to true and at least one variable is set to false?
We give one reduction showing that TTR is NP-hard even if the input tree $G$ has constant diameter and we give a second reduction showing that TTR is NP-hard even if $G$ has constant maximum degree. Both reductions as based on similar ideas. Before we give formal descriptions, we present several properties of labelings that we will exploit in the reductions.
The first property establishes a relation between the duration of a fastest temporal path from $s$ to $t$ and the duration of a fastest temporal path from $t$ back to $s$.
Lemma 3.2.
Let $G$ be a tree and $(G,\lambda)$ be a $\Delta$-periodic temporal graph. Let $P_{s,t}=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$ be a fastest temporal path in $(G,\lambda)$ from $s=v_{0}$ to $t=v_{k}$. Let $P_{t,s}$ be a fastest temporal path in $(G,\lambda)$ from $t$ to $s$. Then we have
$d(P_{s,t})+d(P_{t,s})\geq(k-1)\cdot\Delta+2.$
Furthermore, if for all $i\in[k-1]$ we have $\lambda(\{v_{i-1},v_{i}\})\neq\lambda(\{v_{i},v_{i+1}\})$, then we have
$d(P_{s,t})+d(P_{t,s})=(k-1)\cdot\Delta+2.$
Proof 3.3.
This follows directly from Lemma 2.2 and Section 2.
From Lemma 3.2, we immediately get the following.
Corollary 3.4.
Let $(G,D,\Delta)$ be a yes-instance of TTR. Let $P_{s,t}=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$ with $k>1$ be a fastest temporal path in $(G,\lambda)$ from $s=v_{0}$ to $t=v_{k}$. Assume that $D_{s,t}+D_{t,s}=(k-1)\cdot\Delta+2$. Then for every solution $\lambda$, we have
$\lambda(\{v_{0},v_{1}\})=\lambda(\{v_{k-1},v_{k}\})-D_{s,t}+i\cdot\Delta+1,$
for some (uniquely determined) $i\geq 0$.
Intuitively, Corollary 3.4 allows us to use the matrix $D$ “force” certain labels of certain pairs of edges to be a function of each other. In particular, we can force the labels of certain pairs of edges to be the same. We will heavily exploit this in the gadgets of the reduction.
The next property provides some insight into the maximum duration that any fastest temporal path in a $\Delta$-periodic temporal graph can have.
Lemma 3.5.
Let $(G,\lambda)$ be a $\Delta$-periodic temporal graph. Let $P=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$ be a fastest temporal path in $(G,\lambda)$. Then we have $d(P)\leq(k-1)\cdot\Delta+1$. Furthermore, if $d(P)=(k-1)\cdot\Delta+1$, then we have for all $i\in[k-1]$ that $\lambda(\{v_{i-1},v_{i}\})=\lambda(\{v_{i},v_{i+1}\})$.
Proof 3.6.
This follows directly from Lemma 2.2 and Section 2.
Intuitively, Lemma 3.5 tells us which values in $D$ essentially do not put any constraints on a solution labeling. More formally, given a graph $G$ and two vertices $s,t$ of distance $k$ in $G$, then we say that the upper bound $D_{s,t}$ is trivial if $D_{s,t}=(k-1)\cdot\Delta+1$.
Next, we present variable gadgets and clause gadgets that we use in the reductions. The above properties will help us to prove that the gadgets have the desired functionality. From now on, assume that we are given an instance $(X,C)$ of MonNAE3SAT. We set $\Delta=5$.
For each variable $x\in X$ we create a variable gadget $(G_{x},D)$, where $G_{x}=(V_{x},E_{x})$ with $V_{x}=\{w_{x},v_{x,1},v_{x,2}\}$ and $E_{x}=\{\{w_{x},v_{x,1}\},\{v_{x,1},v_{x,2}\}\}$. We set $D_{w_{x},v_{x,2}}=3$ and make all other upper bounds trivial.
For each clause $c\in C$ we create a clause gadget $(G_{c},D)$, where $G_{c}=(V_{c},E_{C})$ with
  • $V_{c}=\{w_{c},u_{c,1},u_{c,2},u_{c,3},u_{c,4},u_{c,5},u_{c,6}\}$, and
  • $E_{c}=\{\{w_{c},u_{c,1}\},\{u_{c,1},u_{c,2}\},\{u_{c,1},u_{c,3}\},\{u_{c,1},u_{c,4}\},\{u_{c,4},u_{c,5}\},\{u_{c,4},u_{c,6}\}\}$.
Furthermore, we set
  • $D_{u_{c,2},u_{c,4}}=4$ and $D_{u_{c,4},u_{c,2}}=3$,
  • $D_{u_{c,5},u_{c,6}}=4$ and $D_{u_{c,6},u_{c,5}}=3$,
  • $D_{u_{c,3},u_{c,5}}=6$ and $D_{u_{c,5},u_{c,3}}=6$,
  • $D_{w_{c},u_{c,6}}=10$,
and make all other upper bounds trivial.
A visualization of the variable gadget and the clause gadget is given in Figure 2.
Figure 2: Visualization of the variable gadget (left) and the clause gadget (right). Informally, the label on the bold edge of the variable gadget models whether the variable is set to true or false. The three bold edges in the clause gadget correspond to the three variables of the clause. Informally, the clause gadget forbids that all three of those edges obtain the same label, which corresponds to the clause being satisfied if not all three variables are set to the same Boolean value.
For the two hardness results, we arrange the variable and clause gadgets in different ways. One way produces a tree with constant treewidth, and a second way produces a tree with constant maximum degree. The main idea in both cases is the following. Assume the three variables in each clause are ordered in an fixed but arbitrary way.
  • We ensure that all edges $\{w_{x},v_{x,1}\}$ of all variable gadgets $G_{x}$ need to receive the same label. For the sake of creating an intuition, assume this label is one.
  • Then the only non-trivial duration upper bound in the variable gadget $G_{x}$ forces the label on the edge $\{v_{x,1},v_{x_{2}}\}$ of each variable gadget to be either two or three. We interpret choosing label two as setting the variable to true, and choosing label three as setting the variable to false.
  • In each clause gadget $G_{c}$, we associate edge $\{w_{c},u_{c,1}\}$ with the first variable of the clause, say $x$, and use the matrix $D$ to force the label on edge $\{w_{c},u_{c,1}\}$ the same as the label on edge $\{v_{x,1},v_{x_{2}}\}$. Similarly, we associate edge $\{u_{c,1},u_{c,4}\}$ with the second variable $y$ of the clause, and we associate edge $\{u_{c,4},u_{c,6}\}$ with the third variable $z$ of the clause. We use the matrix $D$ to force the label on edge $\{u_{c,1},u_{c,4}\}$ the same as the label on edge $\{v_{y,1},v_{y_{2}}\}$, and to force the label on edge $\{u_{c,4},u_{c,6}\}$ the same as the label on edge $\{v_{z,1},v_{z_{2}}\}$.
  • Then the only non-trivial duration upper bound in the clause gadget $G_{c}$ forces forbids that all three edges $\{w_{c},u_{c,1}\}$, $\{u_{c,1},u_{c,4}\}$, and $\{u_{c,4},u_{c,6}\}$ receive the same label. Intuitively, this ensures that at least one variable in the clause needs to be set to true, and at least one variable in the clause needs to be set to false.
Now we formally describe how to create an instance $(G^{(1)},D^{(1)},\Delta)$ of TTR where $G^{(1)}$ is a tree with constant diameter. We start with two vertices $w^{\star}$ and $w^{\star\star}$ that are connected by an edge $\{w^{\star},w^{\star\star}\}$. Now for each variable $x$, we identify the vertex $w_{x}$ of the variable gadget $G_{x}$ with $w^{\star}$. For each clause $c$, we identify the vertex $w_{c}$ of the clause gadget $G_{c}$ with $w^{\star}$. This finishes the construction of $G^{(1)}$, for an illustration see Figure 3. The following is easy to observe. {observation} The tree $G^{(1)}$ has diameter $6$.
Figure 3: Illustration of $G^{(1)}$ with three variable gadgets for variables $x,y,z$, a clause gadget for clause $c$ and sketched the connection to a second clause gadget for clause $c^{\prime}$. A labeling $\lambda$ is illustrated that would be created (see proof of Lemma 3.7) if clause $c$ contains $x$ as the first, $y$ as the second, and $z$ as the third variable, and $x$ is set to true, $y$ is set to true, and $z$ is set to false.
Now we describe how to construct $D^{(1)}$. First of all, for each pair of vertices from the same gadget, the matrix $D^{(1)}$ adopts the values from the matrix of that gadget. For each variable $x$, we set $D^{(1)}_{v_{x,1},w^{\star\star}}=2$. Let $x$ be a variable that appears in clause $c$. Then we set the following values in $D^{(1)}$.
  • If $x$ is the first variable in $c$, then we set $D^{(1)}_{v_{x,2},u_{c,1}}=6$ and $D^{(1)}_{u_{c,1},v_{x,2}}=6$.
  • If $x$ is the second variable in $c$, then we set $D^{(1)}_{v_{x,2},u_{c,2}}=8$ and $D^{(1)}_{u_{c,2},v_{x,2}}=9$.
  • If $x$ is the third variable in $c$, then we set $D^{(1)}_{v_{x,2},u_{c,3}}=8$ and $D^{(1)}_{u_{c,3},v_{x,2}}=9$.
We set all other entries of $D^{(1)}$ to trivial upper bounds.
This finishes the construction of the instance $(G^{(1)},D^{(1)},\Delta)$. Clearly, we have the following. {observation} The instance $(G^{(1)},D^{(1)},\Delta)$ of TTR can be computed in polynomial time.
In the following we show that $(G^{(1)},D^{(1)},\Delta)$ is a yes-instance of TTR if and only if $(X,C)$ is a yes-instance of MonNAE3SAT. We begin by proving the following direction of the statement.
Lemma 3.7.
If $(X,C)$ is a yes-instance of MonNAE3SAT, then $(G^{(1)},D^{(1)},\Delta)$ is a yes-instance of TTR.
Proof 3.8.
Let $(X,C)$ be a yes-instance of MonNAE3SAT and let $\phi:X\rightarrow\{\text{true},\text{false}\}$ be a solution for $(X,C)$. Then we construct a solution $\lambda$ for $(G^{(1)},D^{(1)},\Delta)$ as follows. See Figure 3 for an illustration of the solution we create.
We set $\lambda(\{w^{\star},w^{\star\star}\})=2$. Now consider the variable gadget for variable $x$. We set $\lambda(\{w^{\star},v_{x,1}\})=1$ and thereby obey the upper bound $D^{(1)}_{v_{x,1},w^{\star\star}}=2$. If $\phi(x)=\text{true}$, when we set $\lambda(\{v_{x,1},v_{x,2}\})=2$, otherwise we set $\lambda(\{v_{x,1},v_{x,2}\})=3$. Note that in both cases the upper bound $D_{w^{\star},v_{x,2}}=3$ is respected (recall the we identified $w_{x}$ with $w^{\star}$). Now we have given labels to all edges corresponding to variable gadgets.
Next, consider clause $c$. Let $x$ be the first variable in $c$. Then we set $\lambda(\{w^{\star},u_{c,1}\})=\lambda(\{v_{x,1},v_{x,2}\})$. It is straightforward to verify that this ensures that both upper bounds $D^{(1)}_{v_{x,2},u_{c,1}}=6$ and $D^{(1)}_{u_{c,1},v_{x,2}}=6$ are respected.
Let $y$ be the second variable in $c$. Then we set $\lambda(\{u_{c,1},u_{c,2}\})=\lambda(\{v_{y,1},v_{y,2}\})+2$ and $\lambda(\{u_{c,1},u_{c,4}\})=\lambda(\{v_{y,1},v_{y,2}\})$. First, note that this means that we set $\lambda(\{u_{c,1},u_{c,2}\})$ either to $4$ or to $5$. Second, we have set $D^{(1)}_{v_{y,2},u_{c,2}}=8$ and $D^{(1)}_{u_{c,2},v_{y,2}}=9$. Third, we have set $D^{(1)}_{u_{c,2},u_{c,4}}=4$ and $D^{(1)}_{u_{c,4},u_{c,2}}=3$. We first argue that the two upper bounds $D^{(1)}_{v_{y,2},u_{c,2}}=8$ and $D^{(1)}_{u_{c,2},v_{y,2}}=9$ are respected. The path in $G^{(1)}$ from $v_{y,2}$ to $u_{c,2}$ follows the following ordered sequence of edges.
$\{v_{y,2},v_{y,1}\},\{v_{y,1},w^{\star}\},\{w^{\star},u_{c,1}\},\{u_{c,1},u_{c,2}\}$
By Lemma 2.2 we know that the duration of a fastest path $P$ from $v_{y,2}$ to $u_{c,2}$ when using the $\Delta$-periodic labeling $\lambda$ on $G^{(1)}$ is
$d(P)=1+\tau_{v_{y,1}}^{v_{y,2},w^{\star}}+\tau_{w^{\star}}^{v_{y,1},u_{c,1}}+\tau_{u_{c,1}}^{w^{\star},u_{c,2}}.$
Furthermore, we know that $\lambda(\{v_{y,1},w^{\star}\})=1$ and $\lambda(\{w^{\star},u_{c,1}\})\in\{2,3\}$. Since $\lambda(\{v_{y,2},v_{y,1}\})>1$ it follows that
  • $\tau_{v_{y,1}}^{v_{y,2},w^{\star}}=6-\lambda(\{v_{y,2},v_{y,1}\})$ (recall that $\Delta=5$),
  • $\tau_{w^{\star}}^{v_{y,1},u_{c,1}}=\lambda(\{w^{\star},u_{c,1}\})-1$, and
  • $\tau_{u_{c,1}}^{w^{\star},u_{c,2}}=\lambda(\{u_{c,1},u_{c,2}\})-\lambda(\{w^{\star},u_{c,1}\})$, since we know that $\lambda(\{u_{c,1},u_{c,2}\})>\lambda(\{w^{\star},u_{c,1}\})$.
It follows that
$d(P)=6-\lambda(\{v_{y,2},v_{y,1}\})+\lambda(\{u_{c,1},u_{c,2}\}).$
Hence, we get that $D^{(1)}_{v_{y,2},u_{c,2}}=8$ imposes the constraint that
$-\lambda(\{v_{y,2},v_{y,1}\})+\lambda(\{u_{c,1},u_{c,2}\})\leq 2.$
Note that setting $\lambda(\{u_{c,1},u_{c,2}\})=\lambda(\{v_{y,1},v_{y,2}\})+2$ fulfills this inequality.
By a symmetric argument, we also get that the constraint imposed by $D^{(1)}_{u_{c,2},v_{y,2}}=9$ is fulfilled. We next show that the upper bounds $D^{(1)}_{u_{c,2},u_{c,4}}=4$ and $D^{(1)}_{u_{c,4},u_{c,2}}=3$ are obeyed. The path in $G^{(1)}$ from $u_{c,2}$ to $u_{c,4}$ first traverses the edge $\{u_{c,2},u_{c,1}\}$ and then the edge $\{u_{c,1},u_{c,4}\}$. By Lemma 2.2 we know that the duration of a fastest path $P$ from $u_{c,2}$ to $u_{c,4}$ when using the $\Delta$-periodic labeling $\lambda$ on $G^{(1)}$ is $d(P)=1+\tau_{u_{c,1}}^{u_{c,2},u_{c,4}}$. We have that $\tau_{u_{c,1}}^{u_{c,2},u_{c,4}}=3$ and $\tau_{u_{c,1}}^{u_{c,4},u_{c,2}}=2$. It follows that both the upper bounds $D^{(1)}_{u_{c,2},u_{c,4}}=4$ and $D^{(1)}_{u_{c,4},u_{c,2}}=3$ are respected.
Let $z$ be the third variable in $c$. Then we set $\lambda(\{u_{c,1},u_{c,4}\})=\lambda(\{u_{c,4},u_{c,5}\})=\lambda(\{v_{z,1},v_{z,2}\})+2$ and $\lambda(\{u_{c,4},u_{c,6}\})=\lambda(\{v_{z,1},v_{z,2}\})$. By an analysis analogous to the one made for the second variable in $c$ it holds that the constraints corresponding to $D^{(1)}_{u_{c,5},u_{c,6}}=4$ and $D^{(1)}_{u_{c,6},u_{c,5}}=3$, $D^{(1)}_{u_{c,3},u_{c,5}}=6$ and $D^{(1)}_{u_{c,5},u_{c,3}}=6$, and $D^{(1)}_{v_{z,2},u_{c,3}}=8$ and $D^{(1)}_{u_{c,3},v_{z,2}}=9$ are fulfilled.
It remains to show that the upper bound imposed by $D^{(1)}_{w^{\star},u_{c,6}}=10$ is obeyed (recall that we identify $w_{c}$ with $w^{\star}$). Assume for contradiction that it is not. Then by Lemma 3.5 we have that
$\lambda(\{w^{\star},u_{c,1}\})=\lambda(\{u_{c,1},u_{c,4}\})=\lambda(\{u_{c,4},u_{c,6}\}).$
However, this can only be the case if variables $x,y,z$ are all set to true or if variables $x,y,z$ are all set to false. In both cases, the clause $c$ is not satisfied, a contradiction. Hence, since all other upper bounds in $D^{(1)}$ are trivial, we can conclude that the constructed labeling $\lambda$ is indeed a solution for $(G^{(1)},D^{(1)},\Delta)$.
We proceed with showing the other direction of the correctness.
Lemma 3.9.
If $(G^{(1)},D^{(1)},\Delta)$ is a yes-instance of TTR, then $(X,C)$ is a yes-instance of MonNAE3SAT.
Proof 3.10.
Let $(G^{(1)},D^{(1)},\Delta)$ be a yes-instance of TTR and let labeling $\lambda$ be solution for the instance. We construct a satisfying assignment for the variables in $X$ as follows.
First, observe that adding a constant (modulo $\Delta$) to all labels does not change any duration of any fastest temporal path in the $\Delta$-periodic temporal graph $(G^{(1)},\lambda)$. Hence, we can assume w.l.o.g. that $\lambda(\{w^{\star},w^{\star\star}\})=2$. Since for all variables $x$ we have that $D^{(1)}_{v_{x,1},w^{\star\star}}=2$, we must also have that $\lambda(\{v_{x,1},w^{\star}\})=1$. Furthermore, since $D^{(1)}_{w^{\star},v_{x,2}}=3$ (recall that we identify $w_{x}$ with $w^{\star}$), we must have that $\lambda(\{v_{x,1},v_{x,2}\})\in\{2,3\}$. If $\lambda(\{v_{x,1},v_{x,2}\})=2$, then we set the variable $x$ to true. Otherwise, we set the variable $x$ to false. We claim that this yields a satisfying assignment for $(X,C)$.
Assume for contradiction that it is not. Then there is a clause $c\in C$ that is not satisfied, that is, either all three variables in the clause are set to true or all three variables are set to false. Assume that $x,y,z$ are the three variables in $c$ and all of them are set to true (the case where all of them are set to false is symmetric). This means that $\lambda(\{v_{x,1},v_{x,2}\})=\lambda(\{v_{y,1},v_{y,2}\})=\lambda(\{v_{z,1},v_{z,2}\})=2$. Assume that w.l.o.g. that $x$ is the first variable of $c$, variable $y$ is second, and $z$ is the third variable of $c$. Note that then we have $D^{(1)}_{v_{x,2},u_{c,1}}=6$ and $D^{(1)}_{u_{c,1},v_{x,2}}=6$, $D^{(1)}_{v_{y,2},u_{c,2}}=8$ and $D^{(1)}_{u_{c,2},v_{y,2}}=9$, and $D^{(1)}_{v_{z,2},u_{c,3}}=8$ and $D^{(1)}_{u_{c,3},v_{z,2}}=9$. Furthermore, the path from $v_{x,2}$ to $u_{c,1}$ in $G^{(1)}$ has length three, and the paths from $v_{y,2}$ to $u_{c,2}$ and from $v_{z,2}$ to $u_{c,3}$ in $G^{(1)}$ each have length four. See Figure 3 for an illustration of $G^{(1)}$. It follows that Corollary 3.4 applies and we can deduce that $\lambda(\{w^{\star},u_{c,1}\})=2$ and $\lambda(\{u_{c,1},u_{c,2}\})=\lambda(\{u_{c,1},u_{c,3}\})=4$. Moreover, we have that $D^{(1)}_{u_{c,2},u_{c,4}}=4$ and $D^{(1)}_{u_{c,4},u_{c,2}}=3$. Since the path from $u_{c,2}$ to $u_{c,4}$ in $G^{(1)}$ has length two, we get by Corollary 3.4 that $\lambda(\{u_{c,1},u_{c,4}\})=2$. Lastly, we have $D^{(1)}_{u_{c,5},u_{c,6}}=4$ and $D^{(1)}_{u_{c,6},u_{c,5}}=3$, and $D^{(1)}_{u_{c,3},u_{c,5}}=6$ and $D^{(1)}_{u_{c,5},u_{c,3}}=6$. The path from $u_{c,3}$ to $u_{c,5}$ in $G^{(1)}$ has length three. By Corollary 3.4 we get that $\lambda(\{u_{c,4},u_{c,5}\})=4$. The path from $u_{c,5}$ to $u_{c,6}$ in $G^{(1)}$ has length two. By Corollary 3.4 we get that $\lambda(\{u_{c,4},u_{c,6}\})=2$. Summarizing, we get that
$\lambda(\{w^{\star},u_{c,1}\})=\lambda(\{u_{c,1},u_{c,4}\})=\lambda(\{u_{c,4},u_{c,6}\})=2.$
This implies that the duration of a fastest path from $w^{\star}$ to $u_{c,6}$ in $(G^{(1)},\lambda)$ is $11$. However, we have $D^{(1)}_{w^{\star},u_{c,6}}=10$ (recall that we identify $w_{c}$ with $w^{\star}$), a contradiction to the assumption that $\lambda$ is a solution for $(G^{(1)},D^{(1)},\Delta)$.
Lemmas 3.7 and 3.9 together with Sections 3 and 3 prove the first part of Theorem 3.1, namely that TTR is NP-hard even if the input tree $G$ has constant diameter. In the remainder, we prove that TTR is NP-hard even if the input tree $G$ has constant maximum degree. We provide a second reduction from MonNAE3SAT that uses the same variable and clause gadgets, but arranges them in a different way. Intuitively, we replace the vertex $w^{\star}$ in $G^{(1)}$ with a long path.
We now formally describe how to create an instance $(G^{(2)},D^{(2)},\Delta)$ of TTR where $G^{(2)}$ is a tree with constant maximum degree. For each variable $x$, we introduce a new vertex $w^{\prime}_{x}$ and connect it to the vertex $w_{x}$ of the variable gadget $G_{x}$, that is, we add edge $\{w_{x},w^{\prime}_{x}\}$. For each clause $c$, we introduce a new vertex $w^{\prime}_{c}$ and connect it to the vertex $w_{c}$ of the clause gadget $G_{c}$, that is, we add edge $\{w_{c},w^{\prime}_{c}\}$. Next, we order the variables in some fixed but arbitrary way and we order the clauses in some fixed but arbitrary way. Let $x$ and $y$ be two variables that are consecutive in the ordering and $y$ is ordered after $x$. Then we connect $w_{x}$ with $w^{\prime}_{y}$, that is, we add edge $\{w_{x},w^{\prime}_{y}\}$. Let $c$ and $c^{\prime}$ be two clauses that are consecutive in the ordering and $c^{\prime}$ is ordered after $c$. Then we connect $w_{c}$ with $w^{\prime}_{c^{\prime}}$, that is, we add edge $\{w_{c},w^{\prime}_{c^{\prime}}\}$. Finally, let $z$ be the variable that is ordered last and let $c^{\prime\prime}$ be the clause that is ordered first. Then we connect $w_{z}$ with $w^{\prime}_{c^{\prime\prime}}$, that is, we add edge $\{w_{z},w^{\prime}_{c^{\prime\prime}}\}$. This finishes the construction of $G^{(2)}$, for an illustration see Figure 4. The following is easy to observe. {observation} The tree $G^{(2)}$ has maximum degree $4$.
Figure 4: Illustration of $G^{(2)}$. Depicted are three variable gadgets for variables $x,y,z$ where $x$ is ordered first, $y$ is ordered second, and $z$ is ordered third. Additionally, a clause gadget for clause $c$ and sketch of a second clause gadget for clause $c^{\prime}$ are depicted, where clause $c^{\prime}$ is ordered (not directly) after clause $c$. A labeling $\lambda$ is illustrated that would be created (see proof of Lemma 3.11) if clause $c$ contains $x$ as the first, $y$ as the second, and $z$ as the third variable, and $x$ is set to true, $y$ is set to true, and $z$ is set to false.
Now we describe how to construct $D^{(2)}$. First of all, for each pair of vertices from the same gadget, the matrix $D^{(2)}$ adopts the values from the matrix of that gadget. For each variable $x$, we set $D^{(2)}_{w^{\prime}_{x},v_{x,1}}=2$. For each pair of variables $x,y$ that are ordered consecutively, where $x$ is ordered before $y$, we set $D^{(2)}_{w^{\prime}_{x},w_{y}}=6$ and $D^{(2)}_{w_{y},w^{\prime}_{x}}=6$.
Let $x$ be a variable that appears in clause $c$. Let $x$ be the $i$th variable in the ordering and let $c$ be the $j$th clause in the ordering. Denote $|X|=n$. Then we set the following values in $D^{(2)}$.
  • If $x$ is the first variable in $c$, then we set
    $D^{(2)}_{v_{x,2},u_{c,1}}=5(n-i+j)+6\text{ and }D^{(2)}_{u_{c,1},v_{x,2}}=5(n-i+j)+6.$
  • If $x$ is the second variable in $c$, then we set
    $D^{(2)}_{v_{x,2},u_{c,2}}=5(n-i+j)+8\text{ and }D^{(2)}_{u_{c,2},v_{x,2}}=5(n-i+j)+9.$
  • If $x$ is the third variable in $c$, then we set
    $D^{(2)}_{v_{x,2},u_{c,3}}=5(n-i+j)+8\text{ and }D^{(2)}_{u_{c,3},v_{x,2}}=5(n-i+j)+9.$
We set all other entries of $D^{(2)}$ to trivial upper bounds.
This finishes the construction of the instance $(G^{(2)},D^{(2)},\Delta)$. Clearly, we have the following. {observation} The instance $(G^{(2)},D^{(2)},\Delta)$ of TTR can be computed in polynomial time.
In the following we show that $(G^{(2)},D^{(2)},\Delta)$ is a yes-instance of TTR if and only if $(X,C)$ is a yes-instance of MonNAE3SAT. This is done in a similar way as in Lemmas 3.7 and 3.9. We begin by proving the following direction of the statement.
Lemma 3.11.
If $(X,C)$ is a yes-instance of MonNAE3SAT, then $(G^{(2)},D^{(2)},\Delta)$ is a yes-instance of TTR.
Proof 3.12.
Let $(X,C)$ be a yes-instance of MonNAE3SAT and let $\phi:X\rightarrow\{\text{true},\text{false}\}$ be a solution for $(X,C)$. Then we construct a solution $\lambda$ for $(G^{(2)},D^{(2)},\Delta)$ as follows. See Figure 4 for an illustration of the solution we create.
Consider the variable gadget for variable $x$. We set $\lambda(\{w^{\prime}_{x},w_{x}\})=5$ and $\lambda(\{w_{x},v_{x,1}\})=1$ and thereby obey the upper bound $D^{(2)}_{w^{\prime}_{x},v_{x,1}}=2$. If $\phi(x)=\text{true}$, when we set $\lambda(\{v_{x,1},v_{x,2}\})=2$, otherwise we set $\lambda(\{v_{x,1},v_{x,2}\})=3$. Note that in both cases the upper bound $D_{w_{x},v_{x,2}}=3$ is respected. Lastly, if $x$ is not the last variable in the ordering, we set $\lambda(\{w_{x},w^{\prime}_{y}\})=4$, where $y$ is the variable directly after $x$ in the ordering, otherwise, we set $\lambda(\{w_{x},w^{\prime}_{c}\})=4$, where $c$ is the first clause in the ordering. Now we have given labels to all edges corresponding to variable gadgets. It is easy to check that for each pair of variables $x,y$ that are ordered consecutively, where $x$ is ordered before $y$, the upper bounds $D^{(2)}_{w^{\prime}_{x},w_{y}}=6$ and $D^{(2)}_{w_{y},w^{\prime}_{x}}=6$ are respected.
Next, consider clause $c$, and let $c$ be the $j$th clause in the ordering. We set $\lambda(\{w^{\prime}_{c},w_{c}\})=5$ and $\lambda(\{w_{c},w^{\prime}_{c^{\prime}}\})=4$, where $c^{\prime}$ is the $(j+1)$st clause (if it exists). Note that this ensures that $\tau^{w_{c^{\prime\prime}},w_{c}}_{w^{\prime}_{c}}=1$ and $\tau^{w^{\prime}_{c},w^{\prime}_{c^{\prime}}}_{w_{c}}=4$, where $c^{\prime\prime}$ is the $(j-1)$st clause (or the last variable). Note that we also have that $\tau^{w_{z},w_{x}}_{w^{\prime}_{x}}=1$ and $\tau^{w^{\prime}_{x},w^{\prime}_{y}}_{w_{x}}=4$ for all variables $x$, where $z$ denotes the variable ordered directly before $x$ (if it exists) and $y$ denotes the variable ordered directly after $x$ (or the first clause).
Let $x$ be the first variable in $c$ and let $x$ be the $i_{x}$th variable in the ordering. Then we set $\lambda(\{w_{c},u_{c,1}\})=\lambda(\{v_{x,1},v_{x,2}\})$. Note that the path from $v_{x,2}$ to $u_{c,t}$ has length $32(n-i+j)+$ 3. A fastest temporal path $P$ from $v_{x,2}$ to $u_{c,t}$ in $(G^{(2)},\lambda)$ traverses $n-i$ vertices $w_{x^{\prime}}$ and $n-i$ vertices $w^{\prime}_{x^{\prime}}$ for some variables $x^{\prime}\neq x$. As observed above travel delays are $4$ and $1$, respectively. Similarly, $P$ traverses $j-1$ vertices $w_{c^{\prime}}$ and $j-1$ vertices $w^{\prime}_{c^{\prime}}$ for some clauses $c^{\prime}\neq c$. Again, as observed above travel delays are $4$ and $1$, respectively. Those travel delays sum up to $5(n-i_{x}+j-1)$. If follows that
$d(P)=1+5(n-i_{x}+j-1)+\tau_{v_{x,1}}^{v_{x,2},w_{x}}+\tau_{w_{x}}^{v_{x,1},w^{\prime}_{x^{\prime}}}+\tau_{w^{\prime}_{c}}^{w_{c^{\prime}},w_{c}}+\tau_{w_{c}}^{w^{\prime}_{c},u_{c,1}},$
where $x^{\prime}$ is the variable ordered directly after $x$ (or the first clause) and $c^{\prime}$ is the clause ordered directly before $c$ (or the last variable). Furthermore, we know that $\lambda(\{v_{x,1},w_{x}\})=1$, $\lambda(\{w_{x},w^{\prime}_{x^{\prime}}\})=4$, $\lambda(\{w^{\prime}_{c},w_{c}\})=5$, and $\lambda(\{w_{c},u_{c,1}\})=\lambda(\{v_{x,1},v_{x,2}\})\in\{2,3\}$. Moreover, we have that $\tau_{w^{\prime}_{c}}^{w_{c^{\prime}},w_{c}}=1$. Since $\lambda(\{v_{x,2},v_{x,1}\})>1$ it follows that
  • $\tau_{v_{x,1}}^{v_{x,2},w_{x}}=6-\lambda(\{v_{x,2},v_{x,1}\})$ (recall that $\Delta=5$),
  • $\tau_{w_{x}}^{v_{x,1},w^{\prime}_{x^{\prime}}}=3$, and
  • $\tau_{w_{c}}^{w^{\prime}_{c},u_{c,1}}=\lambda(\{w_{c},u_{c,1}\})$, since we know that $\lambda(\{w_{c},u_{c,1}\})<5$.
We can conclude that both upper bounds $D^{(2)}_{v_{x,2},u_{c,1}}=5(n-i_{x}+j)+6$ and $D^{(2)}_{u_{c,1},v_{x,2}}=5(n-i_{x}+j)+6$ are respected.
Let $y$ be the second variable in $c$ and let $y$ be the $i_{y}$th variable in the ordering. Then we set $\lambda(\{u_{c,1},u_{c,2}\})=\lambda(\{v_{y,1},v_{y,2}\})+2$ and $\lambda(\{u_{c,1},u_{c,4}\})=\lambda(\{v_{y,1},v_{y,2}\})$. By an analysis analogous to the one made in Lemma 3.7 and for the first variable in $c$ it holds that the constraints corresponding to $D^{(2)}_{u_{c,2},u_{c,4}}=4$ and $D^{(2)}_{u_{c,4},u_{c,2}}=3$, and $D^{(2)}_{v_{y,2},u_{c,2}}=5(n-i_{y}+j)+8$ and $D^{(2)}_{u_{c,2},v_{y,2}}=5(n-i_{y}+j)+9$ are fulfilled.
Let $z$ be the third variable in $c$ and let $z$ be the $i_{z}$th variable in the ordering. Then we set $\lambda(\{u_{c,1},u_{c,4}\})=\lambda(\{u_{c,4},u_{c,5}\})=\lambda(\{v_{z,1},v_{z,2}\})+2$ and $\lambda(\{u_{c,4},u_{c,6}\})=\lambda(\{v_{z,1},v_{z,2}\})$. By an analysis analogous to the one made in Lemma 3.7 and for the first variable in $c$ it holds that the constraints corresponding to $D^{(2)}_{u_{c,5},u_{c,6}}=4$ and $D^{(2)}_{u_{c,6},u_{c,5}}=3$, $D^{(2)}_{u_{c,3},u_{c,5}}=6$ and $D^{(2)}_{u_{c,5},u_{c,3}}=6$, and $D^{(2)}_{v_{z,2},u_{c,3}}=5(n-i_{z}+j)+8$ and $D^{(2)}_{u_{c,3},v_{z,2}}=5(n-i_{z}+j)+9$ are fulfilled.
It remains to show that the upper bound imposed by $D^{(2)}_{w_{c},u_{c,6}}=10$ is obeyed. Assume for contradiction that it is not. Then by Lemma 3.5 we have that
$\lambda(\{w_{c},u_{c,1}\})=\lambda(\{u_{c,1},u_{c,4}\})=\lambda(\{u_{c,4},u_{c,6}\}).$
However, this can only be the case if variables $x,y,z$ are all set to true or if variables $x,y,z$ are all set to false. In both cases, the clause $c$ is not satisfied, a contradiction. Hence, since all other upper bounds in $D^{(2)}$ are trivial, we can conclude that the constructed labeling $\lambda$ is indeed a solution for $(G^{(2)},D^{(2)},\Delta)$.
We proceed with showing the other direction of the correctness.
Lemma 3.13.
If $(G^{(2)},D^{(2)},\Delta)$ is a yes-instance of TTR, then $(X,C)$ is a yes-instance of MonNAE3SAT.
Proof 3.14.
Let $(G^{(2)},D^{(2)},\Delta)$ be a yes-instance of TTR and let labeling $\lambda$ be solution for the instance. We construct a satisfying assignment for the variables in $X$ as follows.
First, observe that adding a constant (modulo $\Delta$) to all labels does not change any duration of any fastest temporal path in the $\Delta$-periodic temporal graph $(G^{(2)},\lambda)$. Hence, we can assume w.l.o.g. that $\lambda(\{w^{\prime}_{x},w_{x}\})=5$, where $x$ is the first variable. Let $y$ be the second variable. We have that $D^{(2)}_{w^{\prime}_{x},w_{y}}=6$ and $D^{(2)}_{w_{y},w^{\prime}_{x}}=6$. It follows by Corollary 3.4 that also $\lambda(\{w^{\prime}_{y},w_{y}\})=5$. By iterating this argument, we obtain that for all variables $x$, we have that $\lambda(\{w^{\prime}_{x},w_{x}\})=5$. Since for all variables $x$ we further have that $D_{w^{\prime}_{x},v_{x,1}}=2$, we must also have that $\lambda(\{v_{x,1},w_{x}\})=1$. Furthermore, since $D_{w_{x},v_{x,2}}=3$, we must have that $\lambda(\{v_{x,1},v_{x,2}\})\in\{2,3\}$. If $\lambda(\{v_{x,1},v_{x,2}\})=2$, then we set the variable $x$ to true. Otherwise, we set the variable $x$ to false. We claim that this yields a satisfying assignment for $(X,C)$.
Assume for contradiction that it is not. Then there is a clause $c\in C$ that is not satisfied, that is, either all three variables in the clause are set to true or all three variables are set to false. Let $c$ be the $j$th clause. Assume that $x,y,z$ are the three variables in $c$ and all of them are set to true (the case where all of them are set to false is symmetric). This means that $\lambda(\{v_{x,1},v_{x,2}\})=\lambda(\{v_{y,1},v_{y,2}\})=\lambda(\{v_{z,1},v_{z,2}\})=2$. Assume that w.l.o.g. that $x$ is the first variable of $c$, variable $y$ is second, and $z$ is the third variable of $c$. Furthermore, let $x$ be the $i_{x}$th variable in the ordering, let $y$ be the $i_{y}$th variable in the odering, and let $z$ be the $i_{z}$th variable in the odering. Note that then we have $D^{(2)}_{v_{x,2},u_{c,1}}=5(n-i_{x}+j)+6$ and $D^{(2)}_{u_{c,1},v_{x,2}}=5(n-i_{x}+j)+6$, $D^{(2)}_{v_{y,2},u_{c,2}}=5(n-i_{y}+j)+8$ and $D^{(2)}_{u_{c,2},v_{y,2}}=5(n-i_{y}+j)+9$, and $D^{(2)}_{v_{z,2},u_{c,3}}=5(n-i_{z}+j)+8$ and $D^{(2)}_{u_{c,3},v_{z,2}}=5(n-i_{z}+j)+9$. Furthermore, the path from $v_{x,2}$ to $u_{c,1}$ in $G^{(2)}$ has length $32(n-i_{x}+j)+$ 3, the path from $v_{y,2}$ to $u_{c,2}$ in $G^{(2)}$ has length $42(n-i_{y}+j)+$ 4, and the path from $v_{z,2}$ to $u_{c,3}$ in $G^{(2)}$ has length $42(n-i_{z}+j)+$ 4. See Figure 4 for an illustration of $G^{(2)}$. It follows that Corollary 3.4 applies and we can deduce that $\lambda(\{w_{c},u_{c,1}\})=2$ and $\lambda(\{u_{c,1},u_{c,2}\})=\lambda(\{u_{c,1},u_{c,3}\})=4$. Moreover, we have that $D^{(2)}_{u_{c,2},u_{c,4}}=4$ and $D^{(2)}_{u_{c,4},u_{c,2}}=3$. Since the path from $u_{c,2}$ to $u_{c,4}$ in $G^{(2)}$ has length two, we get by Corollary 3.4 that $\lambda(\{u_{c,1},u_{c,4}\})=2$. Lastly, we have $D^{(2)}_{u_{c,5},u_{c,6}}=4$ and $D^{(2)}_{u_{c,6},u_{c,5}}=3$, and $D^{(2)}_{u_{c,3},u_{c,5}}=6$ and $D^{(2)}_{u_{c,5},u_{c,3}}=6$. The path from $u_{c,3}$ to $u_{c,5}$ in $G^{(2)}$ has length three. By Corollary 3.4 we get that $\lambda(\{u_{c,4},u_{c,5}\})=4$. The path from $u_{c,5}$ to $u_{c,6}$ in $G^{(2)}$ has length two. By Corollary 3.4 we get that $\lambda(\{u_{c,4},u_{c,6}\})=2$. Summarizing, we get that
$\lambda(\{w_{c},u_{c,1}\})=\lambda(\{u_{c,1},u_{c,4}\})=\lambda(\{u_{c,4},u_{c,6}\})=2.$
This implies that the duration of a fastest path from $w_{c}$ to $u_{c,6}$ in $(G^{(2)},\lambda)$ is $11$. However, we have $D^{(2)}_{w_{c},u_{c,6}}=10$, a contradiction to the assumption that $\lambda$ is a solution for $(G^{(2)},D^{(2)},\Delta)$.
Now we have all ingredients to prove Theorem 3.1.
Proof 3.15 (Proof of Theorem 3.1).
Recall that we set $\Delta=5$. Lemmas 3.7 and 3.9 together with Sections 3 and 3 prove the first part of Theorem 3.1, namely that TTR is NP-hard even if the input tree $G$ has constant diameter. Lemmas 3.11 and 3.13 together with Sections 3 and 3 prove the second part of Theorem 3.1, namely that TTR is NP-hard even if the input tree $G$ has constant maximum degree.
4 FPT-algorithm for TTR parameterized by the number of leaves
Our hardness results from the previous section (Theorem 3.1) exclude FPT-algorithms for a variety of parameters. The parameter we consider here is the number of leaves of the tree, which is larger than the maximum degree of the tree and incomparable to the diameter of the tree.
Theorem 4.1.
TTR is fixed-parameter tractable when parameterized by the number $\ell$ of leaves of the input tree $G$.
To show Theorem 4.1 we present a Turing-reduction from TTR to Mixed Integer Linear Program (MILP).
Mixed Integer Linear Program (MILP)
Input: A vector $x$ of $n$ variables of which some are considered integer variables, a constraint matrix $A\in\mathbb{R}^{m\times n}$, and two vectors $b\in\mathbb{R}^{m}$, $c\in\mathbb{R}^{n}$. Task: Compute an assignment to the variables (if one exists) such that all integer variables are set to integer values, $Ax\leq b$, $x\geq 0$, and $c^{\intercal}x$ is maximized.
Given an instance $(G,D,\Delta)$ of TTR, we will produce several MILP instances. We will prove that $(G,D,\Delta)$ is a yes-instance if and only if at least one of the MILP instances admits a feasible solution. The number of produced instances is upper-bounded by a function of the number $\ell$ of leaves in $G$. Each of the produced MILP instances will have a small number of integer values. More precisely, the number of integer values will be upper-bounded by a function of the number $\ell$ of leaves in $G$. This will allow us to upper-bound the running time necessary to solve the MILP instances using the following known result.
Theorem 4.2 ([48, 20]).
MILP is fixed-parameter tractable when parameterized by the number of integer variables.
Furthermore, we build our MILP formulations in a specific way that ensures that there always exist optimal solutions where all variables are set to integer values. Informally, we ensure that the constraint matrix for the rational variables is totally unimodular. This allows us to use the following result.
Lemma 4.3 ([16]).
Let the following be an MILP.
$\max c^{\intercal}x\text{ subject to }Ax\leq b,x\geq 0.$
Let $x=(x_{\text{int}}\ x_{\text{frac}})^{\intercal}$, where $x_{\text{int}}$ (resp. $x_{\text{frac}}$) denote the integer (resp. fractional) variables of the MILP. Let $A=(A_{\text{int}}\ A_{\text{frac}})$ where $A_{\text{int}}$ are the first $|x_{\text{int}}|$ columns of $A$, that is, the coefficients of the integer variables, and $A_{\text{frac}}$ are the remaining columns, that is, the coefficients of the fractional variables. If $A_{\text{frac}}$ is totally unimodular, then there exists an optimal solution to the MILP where all variables are set to integer values.
The main idea for the reduction is the following. Given an instance $(G,D,\Delta)$ of TTR, we create variables for the travel delays at the vertices of $G$. A labelling can easily be computed from the travel delays. We use the constraints to ensure that the durations of the fastest paths respect the upper bounds given in $D$.
Before we start the description of the reduction, we make a straightforward observation on trees. {observation} Let $G$ be a tree with $\ell$ leaves. Let $V_{>2}$ be the set of vertices in $G$ with degree larger than two. We have that $|V_{>2}|\leq\ell$. Furthermore, we have for ever $v\in V_{>2}$ that the degree of $v$ is at most $\ell$.
Furthermore, we show that there always exist solutions to yes-instances of TTR, where the two edges incident with degree two vertices have different labels.
Lemma 4.4.
Let $(G,D,\Delta)$ be a yes-instance of TTR. Let $V_{2}$ be the set of vertices in $G$ with degree two. Then there exists a solution $\lambda$ for $(G,D,\Delta)$ such that for all $v\in V_{2}$ we have the following.
Let $u,w$ denote the two neighbors of $v$ in $G$. We have that $\lambda(\{u,v\})\neq\lambda(\{v,w\})$.
Proof 4.5.
Assume there is a solution $\lambda$ for $(G,D,\Delta)$ such that for some degree two vertex $v$ in $G$ with neighbors $u,w$ we have that $\lambda(\{u,v\})=\lambda(\{v,w\})$. Then we modify $\lambda$ as follows. Let $G_{u}$ be the subtree of $G$ rooted at $v$ that contains $u$. For each edge $e$ in $G_{u}$, if $\lambda(e)=\Delta$ the we set $\lambda(e)=1$ and otherwise we increase $\lambda(e)$ by one. By Definition 2.1 we have that this leaves all travel delays except $\tau_{v}^{u,w}$ and $\tau_{v}^{w,u}$ unchanged and it decreases $\tau_{v}^{u,w}$ and $\tau_{v}^{w,u}$ each by at least one. It follows no duration of a fastest temporal path increases. By iterating this procedure, we can obtain a solution $\lambda^{\prime}$ with the desired property.
From now on assume that we are given an instance $(G,D,\Delta)$ of TTR. Assume the vertices in $G$ are ordered in an arbitrary but fixed way. Each MILP instance we create will use the same set of variables. For each vertex $v$ in $G$, we create the following variables:
  • If $v$ has degree two, we create a fractional variable $x_{v}$. This variable will correspond to the travel delay $\tau_{v}^{u,w}$, where $u,w$ are the two neighbors of $v$ in $G$ and $u$ is ordered before $w$.
  • If $v$ has degree larger than two, then for every pair $u,w$ of neighbors of $v$ in $G$, we create an integer variable $y_{v}^{u,w}$. This variable will correspond to the travel delay $\tau_{v}^{u,w}$.
    Furthermore, we create an integer variable $z_{e}$ for every edge $e$ that is incident with $v$. This variable will correspond to the label $\lambda(e)$ of $e$.
With the help of Section 4 we can upper-bound the number of created integer variables as follows. {observation} The number of integer variables in each MILP instance is in $O(\ell^{3})$.
Now we consider all possibilities of how the labels of edges incident with vertices of degree larger than two can relate to each other. Formally, let $v$ be a vertex in $G$ that has degree larger than two and let $E_{v}$ be the set of edges incident with $v$. By Section 4 we have that $|E_{v}|\leq\ell$. Any labeling $\lambda$ partitions $E_{v}$ into at most $\min(\ell,\Delta)$ sets of edges with equal labels, and the values of the labels define an ordering of those sets. We call a partitioning of $E_{v}$ into at most $\min(\ell,\Delta)$ sets together with the ordering for those sets a label configuration for $v$. A set of label configurations for all vertices $v$ in $G$ that have degree larger than two is called a global label configuration. With the help of Section 4 we get the following. {observation} There are $O(\ell^{\ell^{2}})$ global label configurations.
For each global label configuration, we create an MILP instance. From now on fix a global label configuration $\sigma$. We describe how to construct an MILP instance $I_{\sigma}$. Consider a vertex $v$ in $G$ that has degree larger than two. Let $u,w$ be a pair of neighbors of $v$ in $G$. Let $e=\{u,v\}$ and $e^{\prime}=\{v,w\}$. If according to the global label configuration, we have that $e$ and $e^{\prime}$ are in different parts of $E_{v}$ and the part of $e$ is ordered before the part of $e^{\prime}$ (that is, $e$ has a smaller label than $e^{\prime}$), then we add the following constraint.
$y_{v}^{u,w}=z_{e^{\prime}}-z_{e}$ (1)
If the part of $e$ is ordered after the part of $e^{\prime}$ (that is, $e$ has a larger label than $e^{\prime}$), then we add the following constraint.
$y_{v}^{u,w}=z_{e^{\prime}}-z_{e}+\Delta$ (2)
In both above cases, we additionally add the following constraint.
$11\leq y_{v}^{u,w}\leq\Delta-$ 1 (3)
If according to the global label configuration, we have that $e$ and $e^{\prime}$ are in the same part of $E_{v}$ (that is, they have the same label), then we add the following constraints.
$z_{e}=z_{e^{\prime}}$ (4)
$y_{v}^{u,w}=\Delta$ (5)
Furthermore, in any of the above cases, we additionally add the following constraints.
$1\leq z_{e}\leq\Delta$ (6)
$1\leq z_{e^{\prime}}\leq\Delta$ (7)
Now we consider all vertex pairs $s,t$ in $G$ that have distance at least two in $G$, that is, the (unique) path from $s$ to $t$ in $G$ has at least one internal vertex. Let $P_{s,t}$ denote the path from $s$ to $t$ in $G$. Then $V^{(s,t)}_{\text{int}}=V(P_{s,t})\setminus\{s,t\}$ is the set of internal vertices of $P_{s,t}$. We partition $V^{(s,t)}_{\text{int}}$ into two sets. Let $V^{(s,t)}_{\text{int},\text{deg}2}\subseteq V^{(s,t)}_{\text{int}}$ be the internal vertices of $P_{s,t}$ with degree two and let $V^{(s,t)}_{\text{int},\text{deg}>2}\subseteq V^{(s,t)}_{\text{int}}$ be the internal vertices with degree larger than two. Now we further partition the two sets of vertices.
We define $V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}\subseteq V^{(s,t)}_{\text{int},\text{deg}2}$ such that for all $v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}$, the neighbor of $v$ that is closer to $s$ is ordered before the neighbor of $v$ that is closer to $t$. Analogously, we define $V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}\subseteq V^{(s,t)}_{\text{int},\text{deg}2}$ such that for all $v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$, the neighbor of $v$ that is closer to $t$ is ordered before the neighbor of $v$ that is closer to $s$.
For the vertices in $V^{(s,t)}_{\text{int},\text{deg}>2}$ we additionally define the following. For $v\in V^{(s,t)}_{\text{int},\text{deg}>2}$ let $v^{\prime}$ denote the neighbor of $v$ that is on the path from $v$ to $s$ and let $v^{\prime\prime}$ denote the neighbor of $v$ that is on the path from $v$ to $t$. We define $u^{(s,t)}(v)=v^{\prime}$ and $w^{(s,t)}(v)=v^{\prime\prime}$.
We add the following constraint for the vertex pair $s,t$.
$\sum_{v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}}x_{v}\ +\sum_{v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}}(\Delta-x_{v})\ +\sum_{v\in V^{(s,t)}_{\text{int},\text{deg}>2}}y^{u^{(s,t)}(v),w^{(s,t)}(v)}_{v}\ \leq\ D_{s,t}-1$ (8)
Finally, for all fractional variables $x_{v}$ we add the following constraint.
$11\leq x_{v}\leq\Delta-$ 1 (9)
This finishes the description of the constraints. Since we only want to check whether a feasible solution exists, we arbitrarily choose the vector $c$ for the objective function to be the all-one vector. This finishes the description of the MILP instance $I_{\sigma}$. Clearly, for a given global label configuration, the MILP instance can be constructed in polynomial time. {observation} Given a global label configuration $\sigma$, the MILP instance $I_{\sigma}$ can be computed in polynomial time.
In the following, we prove the correctness of the algorithm. We first show that if the given instance of TTR is a yes-instance, then at least one of the created MILP instances has a feasible solution.
Lemma 4.6.
Given a yes-instance $(G,D,\Delta)$ of TTR, there exists a global label configuration $\sigma$ such that the MILP instance $I_{\sigma}$ admits a feasible solution.
Proof 4.7.
Let $(G,D,\Delta)$ be a yes-instance of TTR. Let $V_{2}$ be the set of vertices in $G$ with degree two. Then by Lemma 4.4 there exists a solution $\lambda$ for $(G,D,\Delta)$ such that for all $v\in V_{2}$ we have the following. Let $u,w$ denote the two neighbors of $v$ in $G$. We have that $\lambda(\{u,v\})\neq\lambda(\{v,w\})$.
Furthermore, let $\sigma$ be the global label configuration defined by the labeling function $\lambda$. We construct a feasible solution for the MILP instance $I_{\sigma}$ as follows.
First, consider all edges $e$ that are incident with vertices of degree larger than two. We set $z_{e}=\lambda(e)$. Note that, since Constraints (1), (2), and (5) are equality constraints, there is a uniquely defined value for each variable $y_{v}^{u,w}$ that we need to assign to that variable in order to fulfill those constraints. Furthermore, since $\lambda$ defines the global label configuration $\sigma$, we also have that the remaining Constraints (3), (4), (6), and (7) are fulfilled by the assignment of the variables. Lastly, it is obvious that all integer variables are assigned integer values.
Now we assign values to the fractional variables. Let $V_{2}$ denote the set of all degree two vertices in $G$. We have a variable $x_{v}$ for each $v\in V_{2}$. Consider some $v\in V_{2}$ and let $u,w$ denote the two neighbors of $v$, where $u$ is ordered before $w$. Then we set
$x_{v}=\begin{cases}\lambda(\{v,w\})-\lambda(\{u,v\}),&\text{if }\lambda(\{v,w\})>\lambda(\{u,v\}),\\ \lambda(\{v,w\})-\lambda(\{u,v\})+\Delta,&\text{otherwise}.\end{cases}$
Recall that due to Lemma 4.4, we have that $\lambda(\{u,v\})\neq\lambda(\{v,w\})$. It follows that Constraint (9) is fulfilled. It remains to show that Constraint (8) is fulfilled.
To this end, consider a vertex pair $s,t$ and let $P=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$ be a fastest temporal $(s,z)$-path in the $\Delta$-periodic temporal graph $(G,\lambda)$. Since $G$ is a tree, we have that $V(P)$ is uniquely determined. In fact, also the order in which the vertices in $V(P)$ are visited by $P$ is uniquely determined. By Lemma 2.2 we have that
$d(P)=1+\sum_{i\in[k-1]}\tau_{v_{i}}^{v_{i-1},v_{i+1}}.$
Furthermore, we have the following.
  • If $v_{i}$ has degree two in $G$ and $v_{i-1}$ is ordered before $v_{i+1}$, then we have $x_{v_{i}}=\tau_{v_{i}}^{v_{i-1},v_{i+1}}$.
  • If $v_{i}$ has degree two in $G$ and $v_{i+1}$ is ordered before $v_{i-1}$, then by Section 2 we have $x_{v_{i}}=\Delta-\tau_{v_{i}}^{v_{i-1},v_{i+1}}$.
  • If $v_{i}$ has degree larger than two in $G$, then we have $y^{v_{i-1},v_{i+1}}_{v_{i}}=\tau_{v_{i}}^{v_{i-1},v_{i+1}}$.
Lastly, since $\lambda$ is a solution for $(G,D,\Delta)$, we have that $d(P)\leq D_{s,t}$. From this, it follows that Constraint (8) is fulfilled. Hence, we have constructed a feasible solution to the MILP instance $I_{\sigma}$.
Next, we show that if one of the produced MILP instances admits a feasible solution, then we are facing a yes-instance of TTR. To this end, we first show that the constraint matrix for the fractional variables of the produced MILP instances is totally unimodular. Then by Lemma 4.3 we have that if one of the MILP instances admits a feasible solution, then there also is a feasible solution for that instance where all fractional variables are set to integer values. To do this, we use a sufficient condition for matrices to be totally unimodular. Precisely, we use that every so-called network matrix is totally unimodular.
Definition 4.8 (network matrix).
Let $T=(V,A)$ be a directed tree, that is, a tree where each arc has an arbitrary orientation and let $A^{\prime}$ be a set of directed arcs on $V$, the same vertex set as $T$. Let $M$ be a matrix with $|A|$ rows and $|A^{\prime}|$ columns. Then $M$ is a network matrix if the following holds. For all $i,j$ let $e_{i}=(a,b)\in A$ and $e^{\prime}_{j}=(s,t)\in A^{\prime}$.
  • $M_{i,j}=1$ if arc $(a,b)$ appears forwards in a path in $T$ from $s$ to $t$.
  • $M_{i,j}=-1$ if arc $(a,b)$ appears backwards in a path in $T$ from $s$ to $t$.
  • $M_{i,j}=0$ if arc $(a,b)$ does not appear in a path in $T$ from $s$ to $t$.
Lemma 4.9 ([61]).
Network matrices are totally unimodular.
Formally, we show that the produced MILP instances have the following property.
Lemma 4.10.
Let $\sigma$ be a global label configuration and let $I_{\sigma}$ be the corresponding MILP instance. Let $M$ be the constraint matrix for the fractional variables, that is, the constraint matrix obtained by treating the integer variables $y_{v}^{u,w}$ and $z_{e}$ in all constraints of $I_{\sigma}$ as arbitrary constants, deleting all constraints that do not involve any fractional variables, and deleting all duplicate rows. We have that $M$ is a network matrix.
Proof 4.11.
First, notice that Constraints (1), (2), (3), (4), (5), (6), and (7) do not involve fractional variables. Hence, from now on we only consider Constraints (8) and (9), where in Constraint (8) we treat the integer variables as constants. Let $M$ be the resulting constraint matrix where also all duplicate rows are deleted. We show that $M$ is a network matrix.
Figure 5: Illustration of the construction of the directed tree $T$ in Lemma 4.10. On the left an example tree $G$, where the vertices are ordered by their indices. In the middle the line graph $L(G)$ of $G$, where the gray vertices form a maximal clique with more than two vertices. On the right the constructed directed tree $T$, where the vertices of the maximal clique are merged into one vertex (gray) and the directed arcs correspond to the degree two vertices of $G$.
To this end, we define a directed tree $T=(V,A)$ (see Definition 4.8) as follows. We give an illustration in Figure 5. Let $L(G)$ be the line graph of the input tree $G$ of the TTR instance. Now we exhaustively merge all maximal cliques in $L(G)$ with more than two vertices to a single vertex. Let this graph be called $G^{\prime}$. It is easy to see that $G^{\prime}$ is a tree, and every edge in $G^{\prime}$ corresponds to a degree two vertex in $G$. Furthermore, we have that every vertex $v$ in $G^{\prime}$ that has degree larger than two corresponds to a subtree $G_{v}$ (which may be the degenerate tree consisting only of one vertex) in $G$ such that all vertices in $V(G_{v})$ have degree larger than two in $G$. Now we transform $G^{\prime}$ into $T$ by replacing each edge of $G^{\prime}$ with an oriented arc. Consider a degree two vertex $v$ of $G$ and the corresponding edge $e_{v}$ in $G^{\prime}$. Let $u$ and $w$ be the two neighbors of $v$ in $G$, where $u$ is ordered before $w$. We distinguish four cases.
  • Vertex $u$ has degree two in $G$. Let $e_{u}$ be the edge in $G^{\prime}$ corresponding to $u$. Then we orient $e_{v}$ such that the resulting arc points away from the common endpoint of $e_{v}$ and $e_{u}$.
  • The degree of vertex $u$ does not equal two and $w$ has degree two in $G$. Let $e_{w}$ be the edge in $G^{\prime}$ corresponding to $w$. Then we orient $e_{v}$ such that the resulting arc points towards the common endpoint of $e_{v}$ and $e_{w}$.
  • Neither $u$ nor $w$ have degree two. Then we can assume that at least one of $u$ and $w$ has degree larger than two, otherwise $G$ only has three vertices and we can solve the instance in constant time.
    • Vertex $u$ has degree larger than two in $G$. Then let $G_{u}$ denote the maximal subtree of $G$ that contains $u$ such that all vertices in $V(G_{u})$ have degree larger than two in $G$. Then there is a vertex $u^{\prime}$ in $G^{\prime}$ corresponding to $G_{u}$ that has degree larger than two and is incident with $e_{v}$. We orient $e_{v}$ such that the resulting arc points away from $u^{\prime}$.
    • Vertex $u$ has degree one and vertex $w$ has degree larger than two in $G$. Then let $G_{w}$ denote the maximal subtree of $G$ that contains $w$ such that all vertices in $V(G_{w})$ have degree larger than two in $G$. Then there is a vertex $w^{\prime}$ in $G^{\prime}$ corresponding to $G_{w}$ that has degree larger than two and is incident with $e_{v}$. We orient $e_{v}$ such that the resulting arc points towards $w^{\prime}$.
This finishes the description of the directed tree $T$.
Next, we make the following observation: Each Constraint (9) produces the same coefficients for the rational variables in the constraint matrix as one of Constraint (8). To see this, note that each Constraint (9) has exactly one non-zero coefficient for one rational variable $x_{v}$, which is either $1$ or $-1$. Variable $x_{v}$ corresponds to a degree two vertex $v$ in $G$. Let $u$ and $w$ be the two neighbors of $v$ in $G$, where $u$ is ordered before $w$. Then, since $u$ and $w$ have distance at least two in $G$, we have a Constraint (8) corresponding to the pair of vertices $s,t$ with $s=u$ and $t=w$. This constraint has exactly one non-zero coefficient for a rational variable, which is $x_{v}$ and the coefficient is $1$. In the case where $s=w$ and $t=u$ we have the same situation, except that the coefficient is $-1$. Hence, since $M$ has no duplicate rows, we can assume that each row of $M$ corresponds to a Constraint (8).
Now consider a row of $M$ corresponding to Constraint (8), which in turn corresponds to vertex pair $s,t$ in $G$, where $s$ and $t$ have distance at least two in $G$. Let $P_{s,t}$ denote the (unique) path from $s$ to $t$ in $G$. Let $V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}\subseteq V(P_{s,t})$ and $V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}\subseteq V(P_{s,t})$ be defined as in the description of Constraint (8). Then we have that $V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}\cup V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$ are all internal vertices of $P_{s,t}$ that have degree two in $G$. Furthermore, in Constraint (8) corresponding to vertex pair $s,t$ we have that the fractional variable $x_{v}$ has a non-zero coefficient if and only if $v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}\cup V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$. More specifically, the coefficient of $x_{v}$ is $1$ if $v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}$ and the coefficient of $x_{v}$ is $-1$ if $v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$.
Furthermore, by construction we have that each vertex in $V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}\cup V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$ corresponds to an arc of $T$. It remains to show that there is a path $P^{\prime}$ in $T$ that visits exactly the arcs corresponding to the vertices in $V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}\cup V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$ such that the arcs corresponding to vertices in $V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}$ appear forwards in the path and the arcs corresponding to the vertices in $V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$ appear backwards in the path. To this end, we order the vertices in $V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}\cup V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$ in the order in which they are visited by $P_{s,t}$, that is, let $V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}\cup V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}=\{v_{1},v_{2},\ldots,v_{k}\}$. Now we show by induction on $k$ that $P^{\prime}$ exists and that its last arc corresponds to $v_{k}$. If $k=1$ then $P^{\prime}_{s,t}$ only consists of the arc $e_{v_{1}}$ corresponding to $v_{1}$. If $v_{1}\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}$, then we consider $P^{\prime}$ to be the path in which $e_{v_{1}}$ forwards. Otherwise, we consider $P^{\prime}$ to be the path in which $e_{v_{1}}$ backwards. Now assume $k>1$. Let $P^{\prime\prime}$ be the path in $T$ that visits the arcs corresponding to vertices $\{v_{1},\ldots,v_{k-1}\}$ such that for all $i\in[k-1]$ the arc corresponding to $v_{i}$ appears forwards if $v_{i}\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}$ and backwards otherwise. Furthermore, we have that $P^{\prime\prime}$ ends with the arc corresponding to $v_{k-1}$. Now we show that we can append the arc corresponding to $v_{k}$ to $P^{\prime\prime}$. We distinguish two cases.
  • In the first case, $v_{k}$ is visited directly after $v_{k-1}$ by $P_{s,t}$. Then the arcs $e_{v_{k-1}}$ and $e_{v_{k}}$ in $T$ share a common endpoint which is a degree two vertex. It follows that we can append the arc $e_{v_{k}}$ to $P^{\prime\prime}$ to create a new path that ends in $e_{v_{k}}$. We still need to show that $e_{v_{k}}$ is appended in the correct direction. Consider the two neighbors of $v_{k}$ in $G$. One of the neighbors is $v_{k-1}$. Let the second neighbor be $w$. Now assume that $v_{k}\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}$. This implies that $v_{k-1}$ is ordered before $w$. In this case the arc $e_{v_{k}}$ in $T$ points away from the common endpoint of $e_{v_{k-1}}$ and $e_{v_{k}}$, and hence it appears forwards in the $P^{\prime\prime}$ appended with $e_{v_{k}}$. The case where $v_{k}\in V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$ is analogous.
  • In the second case, there is a non-empty set of vertices $\hat{V}$ with degree larger than two, that are visited by $P_{s,t}$ between $v_{k-1}$ and $v_{k}$. Let $\hat{G}$ denote the maximal tree in $G$ that only contains vertices that have degree larger than two in $G$ and that contains $\hat{V}$. By construction, there is vertex $\hat{v}$ in $T$ that corresponds to $\hat{G}$ and both $e_{v_{k-1}}$ and $e_{v_{k}}$ in $T$ have $\hat{v}$ as one of their endpoints. First, we argue that we can append $e_{v_{k}}$ to $P^{\prime\prime}$ and thereby create a path in $T$ that ends with $e_{v_{k}}$. To this end, notice that the vertex visited by $P_{s,t}$ before $v_{k-1}$ cannot be part of $\hat{G}$, otherwise $G$ would contain a cycle. It follows that the common endpoint of $e_{v_{k-1}}$ and $e_{v_{k-1}}$ is not $\hat{v}$ and appending $e_{v_{k}}$ to $P^{\prime\prime}$ indeed produces a path in $T$ that ends with $e_{v_{k}}$. It remains to show that $e_{v_{k}}$ is appended in the correct direction. Consider the two neighbors $u$ and $w$ of $v_{k}$ in $G$. We have that one neighbor is contained in $\hat{G}$, assume w.l.o.g. that it is $u$. Then $u$ is visited by $P_{s,t}$ before $v_{k}$. Now assume that $v_{k}\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}$. This implies that $u$ is ordered before $w$. In this case the arc $e_{v_{k}}$ in $T$ points away from $u$, and hence it appears forwards in the $P^{\prime\prime}$ appended with $e_{v_{k}}$. The case where $v_{k}\in V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}$ is analogous.
It follows that for every row in $M$, we can find a path in $T$ such that the conditions in Definition 4.8 are met. We can conclude that $M$ is a network matrix.
Now we are ready to show that if one of the produced MILP instances admits a feasible solution, then we are facing a yes-instance of TTR.
Lemma 4.12.
Given an instance $(G,D,\Delta)$ of TTR, if there exists a global label configuration $\sigma$ such that the MILP instance $I_{\sigma}$ admits a feasible solution, then $(G,D,\Delta)$ is a yes-instance of TTR.
Proof 4.13.
Let $(G,D,\Delta)$ be an instance of TTR. Assume there exists a global label configuration $\sigma$ such that the MILP instance $I_{\sigma}$ admits a feasible solution. By Lemmas 4.3, 4.9 and 4.10 and the fact that totally unimodularity is preserved under row duplication [61] we have that then there exists a feasible solution where all variables are set to integer values. Assume we are given such a solution for $I_{\sigma}$. We construct a solution $\lambda$ for $(G,D,\Delta)$ as follows.
To this end, we make the following observation. Let $u,v,w$ be three vertices in $G$ such that $\{u,v\}$ and $\{v,w\}$ are edges in $G$ and $u$ is ordered before $w$. Then for each fixed label $\lambda(\{u,v\})$ there is a uniquely defined label $\lambda(\{v,w\})$ such that $\tau_{v}^{u,w}=x_{v}$ if $v$ has degree two in $G$, and $\tau_{v}^{u,w}=y_{v}^{u,w}$ if $v$ has degree larger than two in $G$.
This allows us to produce a labeling $\lambda$ with the following procedure. We arbitrarily pick an edge $e$, say the lexicographically smallest one according to the vertex ordering, in $G$ and set $\lambda(e)=1$. Now we iteratively pick an edge $\{u,v\}$ that has not received a label yet and shares a common endpoint $v$ with an edge $\{v,w\}$ in $G$ that has already received a label. By the observation above, the label of $\{v,w\}$ together with the travel delay $\tau_{v}^{u,w}$ uniquely defines the label of $\{u,v\}$, and in turn $\tau_{v}^{u,w}$ is given by $x_{v}$, $y_{v}^{u,w}$, or $y_{v}^{w,u}$, depending on the degree of $v$ and whether $u$ is ordered before $w$ or not. We claim that the described procedure produces a solution $\lambda$ for $(G,D,\Delta)$.
First, we investigate the travel delays in the $\Delta$-periodic temporal graph $(G,\lambda)$. By construction and Section 2, we have for each vertex $v$ that has degree two in $G$, and neighbors $u,w$, where $u$ is ordered before $w$, that $\tau_{v}^{u,w}=x_{v}$ and $\tau_{v}^{w,u}=\Delta-x_{v}$. For each vertex $v$ in $G$ that has degree larger than two, the above described procedure does not per se guarantee that for each pair $u,w$ of neighbors of $v$ that $\tau_{v}^{u,w}=y_{v}^{u,w}$. However, notice that the labeling $\lambda$ is uniquely determined by the label used for the first edge. It follows that there are $\Delta$ different labelings that can be produced by the above described procedure, one for each label that can be put on the first edge. Call them $\lambda=\lambda_{1},\lambda_{2},\ldots,\lambda_{\Delta}$, where for all edges $e$ and for all $1\leq i<j\leq\Delta$ we have $\lambda_{i}(e)=\lambda_{j}(e)+i-j$. In particular, this means that for each vertex $v$ that has degree larger than two and each edge $e$ that is incident with $v$, there is an $i\in[\Delta]$ such that $\lambda_{i}(e)=z_{e}$. In this case, Constraints (1), (2), (3), (4), (5), (6), and (7) ensure that indeed for all edges $e$ that are incident with $v$ we have that $\lambda_{i}(e)=z_{e}$. Furthermore, these constraints ensure that indeed for each pair $u,w$ of neighbors of $v$ we have $\tau_{v}^{u,w}=y_{v}^{u,w}$. Since all labelings $\lambda_{1},\ldots,\lambda_{\Delta}$ produce the same travel delays, we can conclude that for each vertex $v$ in $G$ that has degree larger than two and for each pair $u,w$ of neighbors of $v$ that in $(G,\lambda)$ we have $\tau_{v}^{u,w}=y_{v}^{u,w}$.
Assume for contradiction that $\lambda$ is not a solution for $(G,D,\Delta)$. Then there are two vertices $s,t$ in $G$ such that the duration of a fastest temporal path $P_{s,t}=\left(\left(v_{i-1},v_{i},t_{i}\right)\right)_{i=1}^{k}$ in the $\Delta$-periodic temporal graph $(G,\lambda)$ is larger than $D_{s,t}$. By Lemma 2.2 we have that
$d(P_{s,t})=1+\sum_{i\in[k-1]}\tau_{v_{i}}^{v_{i-1},v_{i+1}}.$
As discussed above, we have the following:
  • If $v_{i}$ has degree two and $v_{i-1}$ is ordered before $v_{i+1}$, then $\tau_{v_{i}}^{v_{i-1},v_{i+1}}=x_{v_{i}}$.
  • If $v_{i}$ has degree two and $v_{i+1}$ is ordered before $v_{i-1}$, then $\tau_{v_{i}}^{v_{i-1},v_{i+1}}=\Delta-x_{v_{i}}$.
  • If $v_{i}$ has degree larger than two, then $\tau_{v_{i}}^{v_{i-1},v_{i+1}}=y_{v_{i}}^{v_{i-1},v_{i+1}}$.
Hence, using the notation introduced for Constraint (8), we can write
$d(P_{s,t})=1+\sum_{v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{f}}}x_{v}\ +\sum_{v\in V^{(s,t)}_{\text{int},\text{deg}2,\text{b}}}(\Delta-x_{v})\ +\sum_{v\in V^{(s,t)}_{\text{int},\text{deg}>2}}y^{u^{(s,t)}(v),w^{(s,t)}(v)}_{v}.$
However, since Constraint (8) is fulfilled for the vertex pair $s,t$, it follows that $d(P_{s,t})\leq D_{s,t}$, a contradiction. This finishes the proof.
Now we have all ingredients to prove Theorem 4.1.
Proof 4.14 (Proof of Theorem 4.1).
We use the following algorithm to solve TTR. Let $(G,D,\Delta)$ be an instance of TTR. For each global label configuration $\sigma$ we create the MILP instance $I_{\sigma}$. If for some $\sigma$ the instance $I_{\sigma}$ admits a feasible solution, we answer “yes”. Otherwise, we answer “no”.
By Lemmas 4.6 and 4.12 this algorithm is correct. It remains to bound the running time. Let $\ell$ be the number of leaves of $G$. By Section 4 we create $O(\ell^{\ell^{2}})$ MILP instance. By Section 4 we can create each MILP instance in polynomial time and by Section 4 each created MILP instance has $O(\ell^{3})$ integer variables. By Theorem 4.2 it follows that we can solve each MILP instance in FPT-time with respect to parameter $\ell$. We can conclude that the overall running time of the algorithm is in $f(\ell)\cdot|(G,D,\Delta)|^{O(1)}$ for some computable function $f$.
5 Conclusion and future work
We have initiated the investigation of the natural temporal tree realization problem TTR and shown that it NP-hard in quite restrictive cases. On the positive side, we provided an FPT-algorithm for the number of leaves in the input tree as a parameter. Our work spawns several natural future research directions. As an immediate open question, we ask whether the number of leaves can be replaced with a smaller parameter, such as the number of vertices with degree larger than two in the input tree. Another related question is whether TTR admits an FPT-algorithm with the vertex cover number of the input tree as a parameter. Furthermore, we leave open whether TTR can be solved in polynomial time if the input tree is a star. A more general future research direction is to allow general input graphs in TTR. Especially interesting would be graphs with bounded maximum leaf number or bounded vertex cover number, since our results do not imply computational hardness for those cases.